The drawing below shows a right-angled triangle. A straight line crosses the triangle parallel to the line z and encloses an angle of α. The lengths x and y of the bottom and top line segments as well as the angle α are given. Find an equation for the length z. 


1
Expert's answer
2021-08-09T16:02:12-0400
"\dfrac{x\cos\alpha}{z}=\dfrac{x\sin \alpha}{x\sin \alpha+y}"

"z=(x\sin \alpha+y)\cot\alpha"